### v1

Abstract
Distributed learning algorithms, such as the ones employed in Federated Learning (FL), require communication compression to reduce the cost of client uploads. The compression methods used in practice are often biased, which require error feedback to achieve convergence when the compression is aggressive. In turn, error feedback requires client-specific control variates, which directly contradicts privacy-preserving principles and requires stateful clients. In this paper, we propose Compressed Aggregate Feedback (CAFe), a novel distributed learning framework that allows highly compressible client updates by exploiting past aggregated updates, and does not require control variates. We consider Distributed Gradient Descent (DGD) as a representative algorithm and provide a theoretical proof of CAFe’s superiority to Distributed Compressed Gradient Descent (DCGD) with biased compression in the non-smooth regime with bounded gradient dissimilarity. Experimental results confirm that CAFe consistently outperforms distributed learning with direct compression and highlight the compressibility of the client updates with CAFe.
Index Terms:
Distributed Learning, Optimization, Federated Learning, Stochastic Gradient Descent, Compression, Error Feedback.
I Introduction
In distributed learning, a central server coordinates the training of a global model using data stored across multiple clients. The general problem formulation is to minimize the sum of client loss functions, which are typically non-convex. We denote the global model as $x\in\mathbb{R}^{d}$, the client loss functions as $f_{n}:\mathbb{R}^{d}\to\mathbb{R}$, and the global loss function as
$f(x)=\frac{1}{N}\sum f_{n}(x),$ (1)
where $N$ is the number of clients. This formulation is common in Federated Learning (FL) [1], a distributed learning paradigm that has gained popularity due to its privacy-preserving properties. In FL, clients train the global model on their local data and send updates to the server, which aggregates them to update the global model. One of the main challenges in FL is the communication cost associated with transmitting model updates from clients to the central server [2]. This cost can be a major bottleneck, especially when the model is large and the number of clients is substantial. To reduce communication costs, researchers have proposed various compression techniques, such as quantization [3, 4], low-rank factorization [5, 6], sparsification [7, 8, 9], and sketching [1], among others.
To obtain convergence guarantees with aggressive compression, theoretical work often relies on unbiased compression, which is not preferred in practice due to its higher computational cost and lower performance [2, 10]. However, biased compression techniques are known to require error feedback to achieve the same convergence guarantees as their unbiased counterparts [10, 11]. Error feedback, also known as error compensation, requires the server to track the state of each client with a control variate. This is a significant drawback for some distributed learning settings, such as FL, where privacy is a key concern. Moreover, in other distributed learning applications, the server may not have the capability to track the state of each client. Also, in massive cross-device FL, clients are most often stateless [2], which makes error feedback infeasible in this setting as clients cannot keep control variates themselves.
Motivated by the above challenges, we propose Compressed Aggregate Feedback (CAFe), a novel distributed learning framework that allows highly compressible client updates without requiring control variates. Our framework leverages the previous aggregated update at the server to help clients compute a more compressible local update. Clients compress the difference between their local update and the previous aggregated update, and the server adds the previous aggregated update when decoding the received messages. This approach is inspired by error feedback, but does not require control variates, making it compatible with existing privacy mechanisms in FL and suitable for stateless clients. The idea of compressing the compensated errors, for example after motion compensation and temporal prediction, is widely used in video coding [12].
The rest of the paper is organized as follows. In Section II, we discuss related work. In Section III, we present the CAFe framework and provide an overview of the proposed approach. In Section IV, we provide a theoretical analysis of CAFe and compare it to Distributed Compressed Gradient Descent (DCGD) with biased compression. In Section V, we present experimental results that demonstrate the effectiveness of CAFe for standard FL benchmarks. Finally, we conclude the paper in Section VI.
II Related Work
Communication compression is a well-studied topic in distributed learning, as was mentioned in the introduction. For distributed learning compression, error feedback is often suggested to improve convergence guarantees [13]. In [11], the authors study the error feedback mechanism for one-bit per coordinate biased compression. For general sparse compressors, it was studied in [14, 15]. For the decentralized setting, [16, 17] proposed variants of error feedback with general compression operators. For asynchronous methods, [18, 19] also showed that a modified error feedback with general compression operators has good convergence guarantees. In the non-convex setting, [20] showed that error feedback can be used in arbitrarily heterogeneous settings, which was later extended to the stochastic and convex settings in [10].
III CAFe overview
To discuss the algorithm design, first, we must cover some compression preliminaries. When clients send a message to the server, they first encode it using a function $E$. The server decodes the received information using a function $D$. We call these functions the encoder and decoder, respectively. For a general compression mechanism, the composition $D(E(x)):=\mathcal{C}\left({x}\right)$ is called a compression operator [14].
Definition 1.
A compression operator is a function $\mathcal{C}:\mathbb{R}^{d}\to\mathbb{R}^{d}$, paired with a positive compression parameter $\omega<1$, such that for any vector $x$,
$\mathbb{E}\left[{\left\lVert\mathcal{C}\left({x}\right)-x\right\rVert}^{2}\right]\leq\omega{\left\lVert x\right\rVert}^{2}.$ (2)
Example 1 (Top-k compression).
The top-k compression operator sets all but the top $k$ elements of a vector in absolute value to zero. The top-k compression operator has parameter $\omega=1-\frac{k}{d}$ [3].
Next, we describe how compression operators are used when minimizing the global loss function from Eq. 1 in a distributed learning setting. The fundamental algorithm for this purpose is Distributed Compressed Gradient Descent (DCGD) — see Algorithm 1. The pseudocode shows how, at each round, the global model is sent to the clients, which train it using gradients computed with local data. Clients then compress these gradients and send them to the server, which averages them to update the global model. This process is repeated for any desired number of rounds.
Algorithm 1 Distributed Compressed Gradient Descent
1: Input: Global model $x$, Rounds $K$, Encoder-Decoder $(E,D)$ pair for compression, learning rate $\gamma$
2: Initialize global model $x^{0}$
3:for round $k$ from 1 to $K$ do
4:      Send $x^{k}$ to all clients
5:     for each client $n$ in parallel do
6:          $y_{n}^{k}\leftarrow x^{k}-\gamma\nabla f_{n}(x^{k})$triangleright Train $x^{k}$ using local data, store the output in $y_{n}^{k}$
7:          $\Delta_{n}^{k}\leftarrow y_{n}^{k}-x^{k}=-\gamma\nabla f_{n}(x^{k})$triangleright Compute local update
8:          Send $E(\Delta_{n}^{k})$ to server $\triangleright$ Upload local update
9:     end for
10:      $q_{n}^{k}\leftarrow D(E(\Delta_{n}^{k}))$triangleright Decode updates
11:      Aggregate client updates in $\Delta_{s}^{k}:=\frac{1}{N}\sum q_{n}^{k}$
12:      Obtain $x^{k+1}:=x^{k}+\Delta_{s}^{k}$.
13:end for
Note that DCGD is a specific instance of the general distributed learning framework, where we have chosen gradient descent as the optimizer for the local models, and equal-weights averaging for the aggregation strategy. We can derive a general strategy by not determining the choice of aggregation strategy for client updates, and the choice of optimizer for on-client training.
Our framework, CAFe, leverages the previous aggregated update $\Delta_{s}^{k-1}$ to help clients compute a more compressible update. Namely, clients will compress the difference between their local update $\Delta_{n}^{k}$ and the previous aggregated update:
$E(\Delta_{n}^{k}-\Delta_{s}^{k-1}).$
On the server side, the server will add the previous aggregated update when decoding the received messages:
$q_{n}^{k}\leftarrow D(E(\Delta_{n}^{k}-\Delta_{s}^{k-1}))+\Delta_{s}^{k-1}.$
The pseudocode for this procedure is described in Algorithm 2, where the novelty with respect to the general distributed learning framework is highlighted in green boxes.
Algorithm 2 CAFe
1: Input: Global model $x$, Rounds $K$, Encoder-Decoder $(E,D)$ pair for compression
2: Initialize global model $x^{0}$
3: Initialize aggregated update $\Delta_{s}^{0}\leftarrow 0$
4:for round $k$ from 1 to $K$ do
5:      Send $x^{k}$ and $\Delta_{s}^{k-1}$ to all clients $\triangleright$ In the stateful version $\Delta_{s}^{k-1}$ may be omitted.
6:     for each client $n$ in parallel do
7:          $y_{n}^{k}\leftarrow\text{Train}(x_{n}^{k})$triangleright Train $x^{k}$ using local data, store the output in $y_{n}^{k}$
8:          $\Delta_{n}^{k}\leftarrow y_{n}^{k}-x^{k}$triangleright Compute local update
9:          Send $E(\Delta_{n}^{k}-\Delta_{s}^{k-1})$ to server $\triangleright$ Upload difference
10:     end for
11:      $q_{n}^{k}\leftarrow D(E(\Delta_{n}^{k}-\Delta_{s}^{k-1}))+\Delta_{s}^{k-1}$triangleright Decode updates
12:      Aggregate client updates in $\Delta_{s}^{k}$
13:      Obtain $x^{k+1}$ using $x^{k}$ and $\Delta_{s}^{k}$.
14:end for
In the single-client setting, observe how the aggregated update at the server is simply the client update, which means that we can analyze it as a control variate. In this case, CAFe is equivalent to the error feedback mechanism in [20]. However, in the multiple-client setting, the aggregated update is a combination of all client updates, which acts as a proxy for the client-specific control variates and requires novel analysis, shown in Section IV.
Observe that if clients have memory, they can retain $x^{k-1}$. In many popular distributed learning algorithms, $x^{k}$ and $x^{k-1}$ determine $\Delta_{s}^{k-1}$, like Distributed Gradient Descent, FedAvg, etc. This means that the CAFe’s server does not need to send $\Delta_{s}^{k-1}$ if clients have memory. Algorithms with momentum [21] can also easily be adapted to our framework.
IV Analysis
We present analysis for CAFe when using Gradient Descent as the optimizer of choice and a compression operator $\mathcal{C}$ with parameter $\omega<1$. In this case, the iterates of CAFe are:
$x^{k+1}$ $=x^{k}+\Delta_{s}^{k},$ (3)
$\Delta_{s}^{k}$ $=\frac{1}{N}\sum_{n}\mathcal{C}\left({\Delta_{n}^{k}-\Delta_{s}^{k-1}}\right)+\Delta_{s}^{k-1}$ (4)
$\Delta_{n}^{k}$ $=-\gamma\nabla f_{n}(x^{k}).$ (5)
We define $e_{n}^{k}=\mathcal{C}\left({\Delta_{n}^{k}-\Delta_{s}^{k-1}}\right)-\left({\Delta_{n}^{k}-\Delta_{s}^{k-1}}\right)$ as the compression error, and $\hat{e}_{n}^{k}=\frac{e_{n}^{k}}{\gamma}$ as the re-scaled compression error. Then, we obtain
$\Delta_{s}^{k}$ $=\frac{1}{N}\sum_{n}\left({\Delta_{n}^{k}+e_{n}^{k}}\right)$ (6)
$\Delta_{n}^{k}$ $=-\gamma(\nabla f_{n}(x^{k})+\hat{e}^{k}_{n}).$ (7)
Furthermore, we define the average re-scaled compression error as $\overline{e}^{k}:=\frac{1}{N}\sum\hat{e}_{n}^{k}$. Combining these equations, we obtain
$x^{k+1}=x^{k}-\gamma\left({\nabla f(x^{k})+\overline{e}^{k}}\right).$ (8)
Observe that if we have a perfect compressor $\mathcal{C}$, that is, the compression error is zero, we recover Distributed Gradient Descent. For ease of notation, we will denote $g^{k}=\nabla f(x^{k})+\overline{e^{k}}$. Therefore, the iterates of CAFe are $x^{k+1}=x^{k}-\gamma g^{k}$.
For our analysis, we will assume the following:
Assumption 1.
The objective function $f$ is $L$-smooth, which implies that it is differentiable, $\nabla f$ is $L$-Lipschitz, and
$f(y)\leq f(x)+\left\langle{\nabla f(x),y-x}\right\rangle+\frac{L}{2}{\left\lVert y-x\right\rVert}^{2}.$ (9)
Also, the objective function $f$ is lower-bounded by $f^{\star}$.
Assumption 2.
The local gradients have bounded dissimilarity, that is, there exists a $B^{2}\geq 1$ such that
$\frac{1}{N}\sum{\left\lVert\nabla f_{n}(x)\right\rVert}^{2}\leq B^{2}{\left\lVert\nabla f(x)\right\rVert}^{2}.$ (10)
This is a standard assumption, see [2, 13].
Given Assumption 1, and using the fact that $-\left\langle{a,b}\right\rangle=\frac{-{\left\lVert a\right\rVert}^{2}-{\left\lVert b\right\rVert}^{2}+{\left\lVert a-b\right\rVert}^{2}}{2}$, we can ensure
$f(x^{k+1})$ $\leq f(x^{k})-\frac{\gamma}{2}{\left\lVert\nabla f(x^{k})\right\rVert}^{2}+\frac{\gamma}{2}{\left\lVert\nabla f(x^{k})-g^{k}\right\rVert}^{2}$
$-\left({\frac{1}{2\gamma}-\frac{L}{2}}\right){\left\lVert x^{k+1}-x^{k}\right\rVert}^{2}.$ (11)
The second term on the RHS represents the compression error ${\left\lVert\overline{e}^{k}\right\rVert}^{2}$, and we will bound it differently for DCGD and CAFe.
Theorem 1.
Given Assumptions 2 and 1, a positive learning rate $\gamma$ such that $\gamma\leq\frac{1}{L}$, and a compression parameter $\omega<1$, DCGD iterating over $K$ iterations satisfies
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$ $\leq\frac{2F_{0}}{\gamma K\left({1-\omega B^{2}}\right)},$ (12)
where $F_{0}=f(x^{0})-f^{\star}$, as long as $1>\omega B^{2}$.
Proof:
The compression error for Algorithm 1 satisfies
$\mathbb{E}\left[{\left\lVert\overline{e}^{k}\right\rVert}^{2}\right]$ $=\frac{1}{N}\sum_{n}\mathbb{E}\left[{\left\lVert\hat{e}_{n}^{k}\right\rVert}^{2}\right]$ (13)
$\leq\omega\frac{1}{N}\sum_{n}{\left\lVert\nabla f_{n}(x^{k})\right\rVert}^{2}$ (14)
$\leq\omega B^{2}{\left\lVert\nabla f(x^{k})\right\rVert}^{2},$ (15)
where we have used Jensen’s inequality in the first step, Eq. 2 for the compression parameter, and Assumption 2 in the last step. If we assume that $\gamma\leq\frac{1}{L}$, we can simplify Eq. 11 to
$\mathbb{E}\left[f(x^{k+1})\right]\leq f(x^{0})-\frac{\gamma}{2}\left({1-\omega B^{2}}\right)\sum_{\ell=0}^{k}\mathbb{E}\left[{\left\lVert\nabla f(x^{\ell})\right\rVert}^{2}\right],$
where we have telescoped the recursion for $k$ iterations. Averaging over $K$ iterations and re-arranging, we obtain the theorem’s statement. ∎
Corollary 1.
Given Assumptions 2 and 1, and a compression parameter $\omega<1$, DCGD iterating over $K$ iterations results in the following upper bound:
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$ $\leq\frac{2LF_{0}}{K\left({1-\omega B^{2}}\right)},$ (16)
where $F_{0}=f(x^{0})-f^{\star}$.
Proof:
Set $\gamma=1/L$ in Theorem 1. ∎
To analyze DGD with CAFe, we need the following preliminary lemmas.
Lemma 2.
Given an $L$-smooth function $f$, and iterations of the form $x^{k+1}=x^{k}-\gamma g^{k}$, we have
$-\left\langle{\nabla f(x^{k+1}),g^{k}}\right\rangle\leq-\left\langle{\nabla f(x^{k}),g^{k}}\right\rangle+\gamma L{\left\lVert g^{k}\right\rVert}^{2}.$ (17)
Proof:
We have $\left\langle{\nabla f(x^{k}),g^{k}}\right\rangle-\left\langle{\nabla f(x^{k+1}),g^{k}}\right\rangle=\left\langle{\nabla f(x^{k+1})-\nabla f(x^{k}),g^{k}}\right\rangle$, and this can be bounded by
$\left\langle{\nabla f(x^{k+1})-\nabla f(x^{k}),g^{k}}\right\rangle$ $\leq\left\lVert\nabla f(x^{k+1})-\nabla f(x^{k})\right\rVert\left\lVert g^{k}\right\rVert$
$\leq\gamma L{\left\lVert g^{k}\right\rVert}^{2},$
where we have used the $L$-smoothness of $f$ in the last step. Re-arranging, we obtain the desired result. ∎
Lemma 3.
Given Assumptions 2 and 1, the compression error for DGD + CAFe satisfies
$\mathbb{E}\left[{\left\lVert\overline{e}^{k+1}\right\rVert}^{2}\right]$ $\leq\omega\mathbb{E}\left[B^{2}{\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}-{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$ (18)
$+\gamma 2\omega L\mathbb{E}\left[{\left\lVert g^{k}\right\rVert}^{2}\right]+\omega{\left\lVert\overline{e}^{k}\right\rVert}^{2}.$
Proof:
$\mathbb{E}\left[{\left\lVert\overline{e}^{k+1}\right\rVert}^{2}\right]$ $\leq\frac{1}{N}\sum_{n}\mathbb{E}\left[{\left\lVert\hat{e}_{n}^{k+1}\right\rVert}^{2}\right]$
$\leq\frac{\omega}{N}\sum_{n}\mathbb{E}\left[{\left\lVert\nabla f_{n}(x^{k+1})-g^{k}\right\rVert}^{2}\right]$
$=\frac{\omega}{N}\sum_{n}\mathbb{E}\left[{\left\lVert\nabla f_{n}(x^{k+1})\pm\nabla f(x^{k+1})-g^{k}\right\rVert}^{2}\right].$
We can bound the obtained sum by
$\omega\mathbb{E}\left[\left({B^{2}-1}\right){\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}+{\left\lVert\nabla f(x^{k+1})-g^{k}\right\rVert}^{2}\right],$
since the interior product term is null and we can bound the sum of square client gradients using Assumption 2. Now, the last term can be bounded using Lemma 2, since
${\left\lVert\nabla f(x^{k+1})-g^{k}\right\rVert}^{2}$ $={\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}+{\left\lVert g^{k}\right\rVert}^{2}$
$\quad-2\left\langle{\nabla f(x^{k+1}),g^{k}}\right\rangle$
$\leq{\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}+{\left\lVert g^{k}\right\rVert}^{2}$
$\quad-2\left\langle{\nabla f(x^{k}),g^{k}}\right\rangle+2\gamma L{\left\lVert g^{k}\right\rVert}^{2}$
$={\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}-{\left\lVert\nabla f(x^{k})\right\rVert}^{2}$
$\quad+{\left\lVert\nabla f(x^{k})-g^{k}\right\rVert}^{2}+2\gamma L{\left\lVert g^{k}\right\rVert}^{2}.$
Plugging this in to the previous expression we obtain the desired result. ∎
Lemma 4.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be an $L$-smooth function with a lower bound $f^{\star}$. Then, for any $x\in\mathbb{R}^{d}$,
$\|\nabla f(x)\|^{2}\leq 2L(f(x)-f^{\star}).$
Proof:
By Assumption 1, for any $y$, we have:
$f(y)$ $\leq f(x)+\left\langle{\nabla f(x),y-x}\right\rangle+\frac{L}{2}\|y-x\|^{2}.$
We choose $y=x-\frac{1}{L}\nabla f(x)$, and obtain
$f\left(y\right)$ $\leq f(x)-\frac{1}{2L}\|\nabla f(x)\|^{2}.$
Since $f(y)\geq f^{\star}$, we re-arrange and obtain the result. ∎
Theorem 2.
Given Assumptions 2 and 1, a positive learning rate $\gamma$ such that
$\gamma\leq\frac{1-\omega}{L\left({1+\omega}\right)},$ (19)
CAFe + DGD iterating over $K$ iterations results in
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$ $\leq\frac{2F_{0}\left({1-\omega}\right)}{\gamma K\left({1-\omega B^{2}}\right)},$ (20)
where $F_{0}=f(x^{0})-f^{\star}$, as long as $1>\omega B^{2}$.
Proof:
Let us denote $\mathbb{E}\left[f(x^{k+1})+\frac{\gamma}{2\left({1-\omega}\right)}{\left\lVert\overline{e}^{k+1}\right\rVert}^{2}\right]:=\Psi^{k+1}$. Then, if we start from Eq. 11, and add the result from Lemma 3 multiplied by $\frac{\gamma}{2\left({1-\omega}\right)}$, we have
$\Psi^{k+1}$ $\leq-\frac{\gamma}{2}\left({1+\frac{\omega}{1-\omega}}\right)\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$
$-\left({\frac{1}{2\gamma}-\frac{L}{2}-\frac{L\omega}{1-\omega}}\right)\mathbb{E}\left[{\left\lVert x^{k+1}-x^{k}\right\rVert}^{2}\right]$
$+\frac{\gamma}{2}\cdot\frac{\omega B^{2}}{1-\omega}\mathbb{E}\left[{\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}\right]+\Psi^{k}.$ (21)
If $\gamma$ satisfies Eq. 19, we can ignore the second term. Unrolling the recursion for $K$ iterations, we obtain
$\Psi^{K}$ $\leq\Psi^{0}-\frac{\gamma}{2}\left({1+\frac{\omega}{1-\omega}}\right)\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$
$+\frac{\gamma}{2}\cdot\frac{\omega B^{2}}{1-\omega}\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k+1})\right\rVert}^{2}\right].$ (22)
Simplifying, and since the compression error is null at zero,
$f(x^{K})$ $\leq f(x^{0})+\frac{\gamma}{2}\cdot\frac{\omega B^{2}}{1-\omega}\mathbb{E}\left[{\left\lVert\nabla f(x^{K})\right\rVert}^{2}\right]$
$-\frac{\gamma}{2}\left({1+\frac{\omega\left({1-B^{2}}\right)}{1-\omega}}\right)\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right].$ (23)
Next, if we use Lemma 4 to bound the $\mathbb{E}\left[{\left\lVert\nabla f(x^{K})\right\rVert}^{2}\right]$ term, note that $\frac{\gamma\omega B^{2}L}{1-\omega}\leq 1$ is always satisfied since $\omega B^{2}<1$ and Eq. 19 imply it. Thus, we obtain
$\frac{\gamma}{2}\left({1+\frac{\omega\left({1-B^{2}}\right)}{1-\omega}}\right)\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]\leq f(x^{0})-f^{\star}.$
Re-arranging, we obtain the desired result. ∎
Corollary 2.
Given Assumptions 2 and 1, and a compression parameter $\omega<1$, CAFe + DGD iterating over $K$ iterations results in the following upper bound
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\left[{\left\lVert\nabla f(x^{k})\right\rVert}^{2}\right]$ $\leq\frac{2LF_{0}\left({1+\omega}\right)}{K\left({1-\omega B^{2}}\right)},$ (24)
where $F_{0}=f(x^{0})-f^{\star}$.
Proof:
Set $\gamma=\frac{1-\omega}{L\left({1+\omega}\right)}$ in Theorem 2. ∎
Observing Theorems 1 and 2, given a choice of learning rate that satisfies both assumptions, CAFe + DGD outperforms DCGD in terms of convergence rate, as it has a $(1-\omega)$ factor in the numerator that DCGD does not. This can be a significant improvement, especially when the compression parameter is close to 1, which is the case for aggressive compression.
If the learning rate is tuned separately for each approach to be the largest possible, the DCGD’s upper bound is smaller than CAFe + DGD’s, as per Corollaries 1 and 2. However, this is a very aggressive choice of learning rate, and in practice, it is unlikely to be chosen. Also, the difference is a factor $\left({1+\omega}\right)<2$, which is negligible in most cases.
V Experimental Results
TABLE I: Experiment setup.
MNIST EMNIST CIFAR-100
Model CONV4 CONV4 ResNet-18
Learning Rate 0.01 0.01 0.1
# classes (non-iid) 4 4 40
FL Rounds 100 200 2000
V-1 Training Setup
We present FL experiments with 10 clients. The selected datasets are MNIST [22], EMNIST [23], and CIFAR-100 [24], and we follow [25] to choose models for the three datasets, which are CONV4, CONV4, and ResNet-18, respectively. The learning rates are tuned based on the model architectures. We present results for both homogeneous and heterogeneous data cases, denoted iid and non-iid, respectively. For the latter, we randomly sample $40\%$ of total classes for each client. We perform one local training epoch with batch size 512 and vary the number of global training rounds for each experiment. Please see Table I for detailed experimental setup. We run each experiment with 3 random seeds and report the means and standard deviations.
TABLE II: Comparison between CAFe and Direct compression under 4 compression methods: Top-k, Top-k + Quantization, Singular Value Decomposition (SVD), and SVD + Quantization under various compression parameter settings.
Top-k
MNIST EMNIST CIFAR-100
k 0.1 0.01 0.001 0.1 0.01 0.001 0.1 0.01 0.001
iid Direct $95.51\pm 0.30$ $92.99\pm 0.53$ $44.99\pm 3.12$ $81.27\pm 7.49$ $77.95\pm 3.11$ $71.63\pm 1.34$ $\mathbf{39.44\pm 0.66}$ $30.60\pm 1.86$ $16.48\pm 0.35$
CAFe $\mathbf{95.90\pm 0.24}$ $\mathbf{95.05\pm 0.34}$ $\mathbf{91.42\pm 1.25}$ $\mathbf{83.54\pm 4.38}$ $\mathbf{80.19\pm 1.76}$ $\mathbf{75.32\pm 1.93}$ $37.82\pm 1.63$ $\mathbf{37.10\pm 1.96}$ $\mathbf{20.54\pm 1.73}$
non-iid Direct $92.18\pm 0.05$ $89.96\pm 0.37$ $82.63\pm 2.86$ $73.54\pm 0.15$ $71.06\pm 0.61$ $60.33\pm 0.82$ $35.99\pm 3.08$ $\mathbf{24.34\pm 1.49}$ $\mathbf{10.17\pm 0.91}$
CAFe $\mathbf{92.73\pm 0.37}$ $\mathbf{91.15\pm 0.66}$ $\mathbf{88.31\pm 1.53}$ $\mathbf{74.56\pm 0.10}$ $\mathbf{72.07\pm 0.91}$ $\mathbf{63.65\pm 0.61}$ $\mathbf{38.26\pm 2.46}$ $23.11\pm 0.53$ $7.27\pm 1.47$
Top-k + Quantization
MNIST EMNIST CIFAR-100
bits 4 5 6 4 5 6 4 5 6
iid Direct $20.94\pm 13.57$ $92.54\pm 1.49$ $95.13\pm 0.32$ $18.65\pm 22.92$ $77.45\pm 0.26$ $80.90\pm 0.97$ $\mathbf{15.79\pm 6.96}$ $32.72\pm 2.81$ $\mathbf{38.04\pm 0.66}$
CAFe $\mathbf{64.49\pm 37.58}$ $\mathbf{94.04\pm 1.48}$ $\mathbf{95.50\pm 0.37}$ $\mathbf{24.41\pm 30.13}$ $\mathbf{81.12\pm 0.83}$ $\mathbf{82.99\pm 0.65}$ $12.12\pm 4.62$ $\mathbf{33.80\pm 5.94}$ $33.59\pm 0.93$
non-iid Direct $\mathbf{11.35\pm 0.00}$ $\mathbf{36.36\pm 35.37}$ $\mathbf{88.79\pm 2.03}$ $63.82\pm 4.34$ $68.38\pm 2.33$ $72.42\pm 0.65$ $\mathbf{17.23\pm 2.79}$ $\mathbf{29.16\pm 1.36}$ $\mathbf{36.76\pm 0.81}$
CAFe $\mathbf{11.35\pm 0.00}$ $35.94\pm 38.78$ $87.04\pm 5.02$ $\mathbf{70.06\pm 1.97}$ $\mathbf{72.10\pm 1.02}$ $\mathbf{73.89\pm 0.81}$ $5.23\pm 3.05$ $25.86\pm 0.77$ $34.78\pm 2.47$
SVD
MNIST EMNIST CIFAR-100
rank 1 2 3 1 2 3 1 2 3
iid Direct $83.54\pm 1.31$ $90.45\pm 2.90$ $92.42\pm 2.82$ $53.38\pm 2.83$ $69.48\pm 1.27$ $72.84\pm 1.28$ $18.00\pm 1.07$ $25.30\pm 0.70$ $30.03\pm 2.74$
CAFe $\mathbf{93.27\pm 0.48}$ $\mathbf{94.88\pm 0.21}$ $\mathbf{95.34\pm 0.23}$ $\mathbf{77.59\pm 0.28}$ $\mathbf{80.81\pm 0.18}$ $\mathbf{81.26\pm 0.49}$ $\mathbf{4139\pm 0.30}$ $\mathbf{43.21\pm 1.11}$ $\mathbf{42.00\pm 0.70}$
non-iid Direct $78.93\pm 2.44$ $88.36\pm 0.71$ $89.92\pm 0.57$ $\mathbf{32.44\pm 2.08}$ $57.18\pm 1.45$ $64.47\pm 1.13$ $7.63\pm 0.68$ $17.19\pm 1.22$ $21.72\pm 1.36$
CAFe $\mathbf{87.00\pm 0.83}$ $\mathbf{91.02\pm 2.58}$ $\mathbf{93.04\pm 0.35}$ $38.57\pm 2.65$ $\mathbf{68.10\pm 0.98}$ $\mathbf{72.23\pm 0.65}$ $\mathbf{11.14\pm 1.03}$ $\mathbf{30.35\pm 1.70}$ $\mathbf{34.54\pm 1.55}$
SVD + Quantization
MNIST EMNIST CIFAR-100
bits 2 3 4 2 3 4 2 3 4
iid Direct $68.94\pm 26.21$ $85.84\pm 0.99$ $83.59\pm 0.71$ $35.18\pm 16.71$ $52.77\pm 3.04$ $51.15\pm 1.54$ $12.78\pm 0.49$ $16.81\pm 1.39$ $18.43\pm 0.17$
CAFe $\mathbf{90.45\pm 0.46}$ $\mathbf{92.77\pm 0.41}$ $\mathbf{93.29\pm 0.42}$ $\mathbf{55.67\pm 5.97}$ $\mathbf{72.43\pm 1.11}$ $\mathbf{77.19\pm 0.36}$ $\mathbf{20.28\pm 2.91}$ $\mathbf{32.57\pm 2.74}$ $\mathbf{38.55\pm 0.92}$
non-iid Direct $\mathbf{66.51\pm 2.65}$ $79.73\pm 0.75$ $79.22\pm 2.09$ $\mathbf{14.31\pm 4.15}$ $30.11\pm 2.71$ $3.73\pm 2.54$ $4.66\pm 0.04$ $7.53\pm 0.48$ $7.78\pm 0.94$
CAFe $63.33\pm 10.19$ $\mathbf{86.33\pm 0.30}$ $\mathbf{86.96\pm 0.37}$ $6.23\pm 5.80$ $\mathbf{41.76\pm 0.65}$ $\mathbf{40.12\pm 1.51}$ $\mathbf{11.41\pm 0.42}$ $\mathbf{12.52\pm 1.08}$ $\mathbf{12.01\pm 1.12}$
(a) MNIST - iid
(b) MNIST - non-iid
(c) EMNIST - iid
(d) EMNIST - non-iid
(e) CIFAR100 - iid
(f) CIFAR100 - non-iid
Figure 1: SVD compression examples for (a) MNIST - iid, (B) MNIST - non-iid, (c) EMNIST - iid, (d) EMNIST - non-iid, (e) CIFAR100 - iid, and (f) CIFAR100 - non-iid. The learning rates have been tuned to be as high as possible while maintaining convergence. Experimentally, these are equal for both methods.
V-2 Compression Setup
We show the effectiveness of CAFe compared with direct compression under 4 biased compression methods: Top-k, Top-k + Quantization, Singular Value Decomposition (SVD) [6], and SVD + Quantization under various compression parameter settings. Specifically, we select $k=\{10\%,1\%,0.1\%\}$ of the total coordinates Top-k compression, top singular vector for SVD, uniform quantization with $bits=\{4,5,6\}$ for Top-k + Quantization, and $bits=\{2,3,4\}$ for SVD + Quantization. For our Quantization experiments, we fix k to $10\%$ for Top-k and rank to $1$ for SVD. Also, we sparsify first and quantize after, as is optimal for FL [26].
V-3 Results
The comparison results between CAFe and direct compression can be found in Table II. Observe that the results align with our theory: CAFe outperforms direct compression in moderate heterogeneity settings (MNIST and EMNIST), while it may suffer when the heterogeneity is higher and compression is very aggressive (CIFAR-100). For Top-k + Quantization compression, we aim to test the lower limit of the choice for the number of bits per coordinate. Notice that when the number of bits is smaller than 5, both CAFe and direct compression result in low performance and large variance, which indicates that it is more suitable to choose $n_{bits}>5$ for this compression method. For SVD compression, CAFe consistently outperforms direct compression by a large margin, regardless of the model architecture and dataset. This is due to SVD’s low compression error, even when using rank 1. With SVD + Quantization, we also test the lower limit and find that it is suitable to choose $n_{bits}>2$.
Since SVD provides a high level of compression with low bitrate, we show the convergence rates by plotting the learning curves using SVD compression in Fig. 1. It can be observed that CAFe achieves not only better performance but faster convergence compared to direct compression.
VI Conclusion
In this paper, we presented CAFe, a novel distributed learning framework that reduces communication costs by leveraging compressible client update messages and sending the previous aggregated update to clients. Our method significantly reduces the upload communication cost for biased compressors used in practice, and we prove analytical guarantees of convergence when optimizing locally with Gradient Descent. Experimental results demonstrate that CAFe outperforms its predecessors when using the same compressors, making it a promising approach for bandwidth-efficient distributed learning.